If both of John's parents are heterozygous for sickle cell anemia and Sarah's parents are AA and Aa what are the chances that John and Sarah's child will have Sickle Cell Anemia? (Sickle Cell Anemia is an autosomal recessive disease)
Answer #1

Sickle cell disease is an autosomal recessive condition. Assume that the gene coding for normal phenotype is A and the gene coding for sickle cell phenotype is a; A is dominant over a.

Given that, John's parents are heterozygous for sickle cell anemia.

Aa* Aa ----> AA (1/4), Aa (1/2), aa (1/4)

The probability of John to be a heterozygous carrier is 2/3. The probability of Sarah to be a carrier is 1/2.

The probability that the Jogn and Sarah will have Sickle cell anemia is = 1/2*2/3*1/2*1/2 = 1/12